# Unveiling the Basin-Like Loss Landscape in Large Language Models

### v2

@@ 1 Introduction
  • Why does fine-tuning with benign data sometimes compromise capabilities acquired during prior alignment (Qi et al., 2023; Du et al., 2024; Mukhoti et al., 2023; Bianchi et al., 2023; Lyu et al., 2024; Hsu et al., 2024; Li et al., 2025; Liu et al., 2024b)?
@@
We posit that these issues can be partially explained by the loss landscape of LLMs (Li et al., 2018). To investigate this, we analyze two complementary perspectives: the most-case landscape, capturing capacity degradation when parameters move along most directions, and the worst-case landscape, reflecting degradation along the worst direction. Our main findings are summarized below.
@@
Figure 1: The most-case loss landscape of different models. Specific benchmarks and visualization details are provided in Sec.˜3.2. As shown, the loss landscape of LLMs resembles a basin, within which models perform nearly identically and outside of which they lose all capabilities.
As shown in Fig.˜1, the most-case landscape exhibits a basin-like structure: models perform nearly identically within the basin, but rapidly lose capabilities once outside (Peng et al., 2024). This basin gradually emerges and expands as model size increases (see Fig.˜7), consistent with recent evidence that LLMs are robust to common parameter perturbations (Men et al., 2024; Gromov et al., 2024; Sun et al., 2023). In particular, pre-training forms a broad “basic capability basin” that confers fundamental language and conversational skills, while subsequent alignment stages carve out narrower “specific capability basins” (e.g., safety, math, coding) in proximity to it. These observations suggest that benign fine-tuning constrained within the basin preserves prior capabilities.
@@
Through above exploratory studies, we construct a smooth model such that the performance degradation along the worst-case direction is theoretically bounded by the size of the most-case basin. This implies that we can derive a theoretical bound on performance degradation for any fine-tuning and input jailbreaking. Together with our conjecture that benign fine-tuning preserves capabilities within the most-case basin, we conclude that enlarging the most-case basin ❶ enhances benign fine-tuning, ❷ mitigates adversarial fine-tuning, and ❸ improves robustness against input jailbreaking. Additionally, we demonstrate that the basins can be readily enlarged during pre-training, likely due to the over-parameterization property of neural networks (Belkin et al., 2019; Allen-Zhu et al., 2019). We hope that our work sheds light on the relationships between loss landscapes, robustness to benign and harmful fine-tuning, jailbreaking, and our proposed theoretical lower bounds and optimization strategies inspire future large-scale studies on pre-training and fine-tuning.
@@ 3.1 Visualization of the Loss Landscape
The loss landscape visualizes the performance change of a neural network w.r.t. parameter perturbations. Formally, let $f_{\bm{\theta}}$ denote a language model with parameters $\bm{\theta}\in\mathbb{R}^{d}$, and let $J_{f,\mathcal{D}}$ represent the benchmark functional on a dataset $\mathcal{D}$, defined as $J_{f,\mathcal{D}}({\bm{\theta}}):=\mathbb{E}_{\bm{x}\in\mathcal{D}}[\mathcal{O}(f_{\bm{\theta}}(\bm{x}))]$, which takes $f_{\bm{\theta}}$ as input and returns a benchmark value on $\mathcal{D}$, characterizing its specific capability. $\mathcal{O}$ is the judgment oracle that takes the output of $f_{\bm{\theta}}$ and returns 0 or 1 based on its correctness/safety. The higher the benchmark value $J_{f,\mathcal{D}}(\bm{\theta})$, the better the performance of the evaluated model $f_{\bm{\theta}}$.
@@
However, there are still several issues with visualizing the loss landscape. First, while lower values of loss indicate better performance, higher values of the benchmark $J_{f,\mathcal{D}}(\bm{\theta})$ are preferable. Moreover, benchmark scores are not directly comparable across tasks (e.g., MMLU scores range from at least 0.25 and rarely exceed 0.8), which can lead to misleading interpretations. To address this, we apply a transformation that flips the benchmark values by subtracting them from one and then normalizes them via min-max normalization. We denote this operation by $\mathcal{T}$, i.e., the visualized value is $\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta})$. The raw benchmark results are reported in Sec.˜D.2.
Besides, directly visualizing a $d$-dimensional landscape is computationally expensive and unintuitive (Li et al., 2018). We follow the common practice (Goodfellow et al., 2014; Im et al., 2016; Smith & Topin, 2017; Li et al., 2018): for a 2-D landscape, we visualize the loss landscape along a specific direction $\bm{\delta}\in\mathbb{R}^{d}$, reducing the problem to visualizing a single-variable function: $L(\alpha)=\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$. For a 3-D landscape, we visualize along two random directions $\bm{\delta}_{1},\bm{\delta}_{2}\in\mathbb{R}^{d}$ using $L(\alpha,\beta)=\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta}_{1}+\beta\bm{\delta}_{2})$.
@@ 3.2 Most-case Loss Landscape
$L(\alpha)=\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta}),\;\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I}).$ (1)
Empirically, we observe that different directions $\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})$ yield nearly identical results. Since $\mathcal{N}(\bm{0},\bm{I})$ represents uniformly random directions, this suggests that most directions produce similar landscapes. Therefore, we designate this as the most-case loss landscape.
General Geometry. As shown in Fig.˜1, the most-case loss landscape for each capability resembles a basin, within which the models perform nearly identically, and outside of which they rapidly lose all capabilities (Peng et al., 2024). In particular, the pre-training stage creates a “basic capability basin” that endows the model with fundamental language comprehension and conversational abilities. Subsequent alignment stages sequentially establish specific capability basins (e.g., safety (Zou et al., 2023), math (Cobbe et al., 2021), coding (Mark Chen, 2021)) near this basic capability basin. More interestingly, this “basin” phenomenon emerges and becomes larger as the model size increases. For models like the Qwen-0.5 model, the loss landscape resembles the 0-1 loss landscape for small models in Li et al. (2018), which is continuous and smooth. When the models become bigger, the basins become more significant and easier to observe. Based on this observation, we argue that as long as subsequent benign fine-tuning remains within the basin of a specific capability, the parameters will remain within this basin and thus will not compromise those capabilities.
@@
Hypothesis Testing. A common concern is whether the loss landscape along several random directions can represent the geometry of a high-dimensional landscape. In fact, if a property holds across several random directions, hypothesis testing enables us to assert, with an arbitrary type-I error, the percentage of directions that satisfy this property. For example, determining the percentage of directions in a $d$-dimensional space that form a literal basin is a hypothesis testing problem aimed at obtaining a statistical lower bound for the expectation of a Bernoulli variable $\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathbb{I}\{\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0\}]$. We use the Clopper-Pearson bound here; see Sec.˜D.4 for detailed configurations and results.
Disclaimer. Note that the basin phenomenon occurs only when using generative-based benchmarks. When using likelihood-based benchmarks, the loss landscape remains smooth and continuous. See Appendix˜C for details. However, the loss landscape using benchmarks remains a valid loss landscape, since all benchmarks used in this paper employ a 0-1 loss on the dataset $\mathcal{D}$, where each sample is assigned a loss of 0/1 for a correct/incorrect response, as shown in Sec.˜3.1. The landscape of 0-1 loss is widely used and studied (Keskar et al., 2017; Li et al., 2018; Garipov et al., 2018).
@@ 3.3 Worst-case Loss Landscape
The constraint $\|\bm{\delta}\|_{2}^{2}=\mathbb{E}[\|\mathcal{N}(\bm{0},\bm{I})\|_{2}^{2}]$ ensures that the perturbation norm matches that used in Figs.˜1 and 2, enabling more direct comparison for each model. Eq.˜2 is solved by optimizing $L(\bm{\theta}+b\bm{\delta})$ using SGD and projecting the norm of $\bm{\delta}$ to unity at each step (Madry et al., 2018).
@@ 3.4 SFT-case Loss Landscape
Results. As shown in Fig.˜3(a), the loss landscape along the benign fine-tuning direction resembles the most-case landscape. It preserves safety within the most-case basin and loses capability when moving outside this basin. In Fig.˜3(b), when there is a distributional gap between the SFT dataset and the original dataset, the loss landscape becomes narrower and sharper, indicating that the fine-tuning direction does not align with the most-case directions. In Fig.˜3(c), when fine-tuning on the adversarial dataset, the model rapidly loses all capabilities, responding only with phrases like “Sure, here is” without providing factual answers to questions. Thus, while some SFT configurations align closely with the most-case direction (e.g., Fig.˜3(a)), others deviate and degrade capabilities more rapidly (e.g., Fig.˜3(b)). This variation clearly depends on the dataset and hyperparameters.
@@ 4 Theoretical Benefits of Basins
To begin, we first formally define the size of the most-case basin. We adopt a soft definition of basins, similar to the flatness definition in Andriushchenko et al. (2023):
Definition 4.1
A model $f_{\bm{\theta}}$ is said to have a $\sigma$-basin on benchmark $J_{\mathcal{D}}$, if its noised version $f_{\bm{\theta}+\bm{\epsilon}}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$, performs nearly the same as the original version $f_{\bm{\theta}}$, i.e.,
$J_{f,\mathcal{D}}(\bm{\theta})-\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\leq\epsilon.$ (3)
Definition 3 is a necessary condition for a model to have a most-case landscape resembling that in Fig.˜1: when $\epsilon\to 0$, it becomes the strict definition of basins, as defined by $\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathbb{I}\{\mathcal{T}\circ J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0\}]$ in Sec.˜3.2 (assuming performance does not increase under noise perturbation). Since our theoretical analysis holds for all $\epsilon$ rather than only $\epsilon=0$, we adopt this definition in this section to provide a theoretical analysis for a more general case.
@@ 4.1 Any Alignment Can Be Bounded by Average-case Alignment
This is achieved through the concept of randomized smoothing (Cohen et al., 2019; Salman et al., 2019; Lee et al., 2019). Since the model performs nearly identically within a $\sigma$-basin, for any input $\bm{x}$, instead of returning $f_{\bm{\theta}}(\bm{x})$, we can sample $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and return $f_{\bm{\theta}+\bm{\epsilon}}(\bm{x})$. Thus, the benchmark value for this model is $J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})$. The following theorem demonstrates that, for any bounded benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$ [^2] Without loss of generality, any benchmark with a bounded output range can be normalized to this interval to obtain a corresponding certified bound., regardless of how sensitive $f$ is to its parameters, the smoothed model $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ is at most $\frac{1}{\sqrt{2\pi}\sigma}$-Lipschitz. Consequently, when $\bm{\theta}_{0}$ is updated to $\bm{\theta}_{sft}$, the benchmark value changes by at most $\frac{1}{\sqrt{2\pi}\sigma}\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}$.
Theorem 4.2
(Weak Law of Randomized Smoothing (Salman et al., 2019)) For any benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, the function $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ is at most $\frac{1}{\sqrt{2\pi}\sigma}$-Lipschitz. Thus, we can bound the performance degradation as:
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\geq\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]-\frac{1}{\sqrt{2\pi}\sigma}\cdot\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}.$ (4)
@@
(Strong Law of Randomized Smoothing (Cohen et al., 2019; Salman et al., 2019; Lee et al., 2019; Chen et al., 2025a)) For any benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, we have:
@@
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}\right),$ (5)
where $\Phi$ is the cumulative distribution function of the standard Gaussian distribution $\mathcal{N}(0,1)$, i.e.,
@@
$\Phi(x)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{x}\exp\left(-\frac{s^{2}}{2}\right)ds.$
@@
The term $\mathbb{E}_{\bm{\epsilon}}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ represents the expected benchmark value when sampling $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and evaluating the benchmark on $f_{\bm{\theta}+\bm{\epsilon}}$. Eq.˜5 provides a guarantee of the expected benchmark value during fine-tuning. In practice, one typically samples a single $\bm{\epsilon}$ and evaluates $f_{\bm{\theta}+\bm{\epsilon}}$, rather than computing the expectation via extensive Monte Carlo sampling. Empirically, sampling a single instance yields results comparable to multiple samples. Theoretically, the variance introduced by sampling can be bounded using the well-known concentration phenomenon, which states that a random variable is unlikely to deviate significantly from its expectation:
Theorem 4.4
@@
(Concentration of Gaussian, adapted from Wainwright (2019)) With probability at least $1-\delta$, we have:
$J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})\geq\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]-L\sigma\sqrt{2\log\frac{1}{\delta}},$ (6)
where $L$ is the Lipschitz constant of $J_{f,\mathcal{D}}$ with respect to $\bm{\theta}$.

### v3

@@ 1 Introduction
  • Why does fine-tuning with benign data sometimes compromise capabilities acquired during prior alignment (Qi et al., 2023; Du et al., 2024; Mukhoti et al., 2023; Bianchi et al., 2023; Lyu et al., 2024; Hsu et al., 2024; Li et al., 2025a; Liu et al., 2024b)?
@@
We posit that these issues can be partially explained by the loss landscape of LLMs (Li et al., 2018). Specifically, we study the loss landscape defined by generative benchmark evaluations (i.e., 0-1 task success) rather than the smooth likelihood surface, as this metric directly captures the stability of model capabilities (see Sec.˜3.1). To investigate this, we analyze two complementary perspectives: the most-case landscape, capturing capacity degradation when parameters move along most directions, and the worst-case landscape, reflecting degradation along the worst direction.
@@
Figure 1: The most-case loss landscape of different models. Specific benchmarks and visualization details are provided in Sec.˜3.2. As shown, the loss landscape of LLMs resembles a basin, within which models perform nearly identically and outside of which they lose all capabilities. The raw data are presented in Table˜1, and the 3D version is shown in Fig.˜6.
As shown in Fig.˜1, the most-case landscape exhibits a basin-like structure: models perform nearly identically within the basin, but rapidly lose capabilities once outside (Peng et al., 2024). This basin gradually emerges and expands as model size increases (see Fig.˜8), consistent with recent evidence that LLMs are robust to common parameter perturbations (Men et al., 2024; Gromov et al., 2024; Sun et al., 2023). In particular, pre-training forms a broad “basic capability basin” that confers fundamental language and conversational skills, while subsequent alignment stages carve out narrower “specific capability basins” (e.g., safety, math, coding) in proximity to it. These observations suggest that benign fine-tuning constrained within the basin preserves prior capabilities.
@@
Through above exploratory studies, we construct a smooth model such that the performance degradation along the worst-case direction is theoretically bounded by the size of the most-case basin. This implies that we can derive a theoretical bound on performance degradation for any fine-tuning and input jailbreaking. Together with our conjecture that benign fine-tuning preserves capabilities within the most-case basin, we conclude that enlarging the most-case basin ❶ enhances benign fine-tuning, ❷ mitigates adversarial fine-tuning, and ❸ improves robustness against input jailbreaking. Experimentally, we demonstrate that (1) the basins can be readily enlarged during pre-training, likely due to the over-parameterization property of neural networks (Belkin et al., 2019; Allen-Zhu et al., 2019); and (2) explicitly optimizing for robustness against Gaussian perturbations effectively mitigates catastrophic forgetting, validating our premise that Gaussian noise serves as a both empirical and theoretical upper bound for benign fine-tuning degradation. We hope that our work sheds light on the relationships between loss landscapes, robustness to benign and harmful fine-tuning, jailbreaking, and our proposed theoretical lower bounds and optimization strategies inspire future large-scale studies on pre-training and fine-tuning.
@@ 3.1 Visualization of the Loss Landscape
The loss landscape visualizes the performance change of a neural network w.r.t. parameter perturbations. Formally, let $f_{\bm{\theta}}$ denote a language model with parameters $\bm{\theta}\in\mathbb{R}^{d}$, and let $\mathcal{S}_{f,\mathcal{D}}$ represent the benchmark score functional on a dataset $\mathcal{D}$, defined as $\mathcal{S}_{f,\mathcal{D}}({\bm{\theta}}):=\mathbb{E}_{\bm{x}\in\mathcal{D}}[\mathcal{O}(f_{\bm{\theta}}(\bm{x}))]$, which takes $f_{\bm{\theta}}$ as input and returns a benchmark value on $\mathcal{D}$, characterizing its specific capability. $\mathcal{O}$ is the judgment oracle that takes the output of $f_{\bm{\theta}}$ and returns 0 or 1 based on its correctness/safety. The higher the benchmark score $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta})$, the better the performance of the evaluated model $f_{\bm{\theta}}$.
@@
However, there are still several issues with visualizing the loss landscape. First, while lower values of loss indicate better performance, higher values of the benchmark $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta})$ are preferable. Moreover, benchmark scores are not directly comparable across tasks (e.g., MMLU scores range from at least 0.25 and rarely exceed 0.8), which can lead to misleading interpretations. To address this, we apply a transformation that flips the benchmark values by subtracting them from one and then normalizes them via min-max normalization. We denote this operation by $\mathcal{T}$, i.e., the visualized value is $\mathcal{T}\circ\mathcal{S}_{f,\mathcal{D}}(\bm{\theta})$. The raw benchmark results are reported in Sec.˜D.2.
Besides, directly visualizing a $d$-dimensional landscape is computationally expensive and unintuitive (Li et al., 2018). We follow the common practice (Goodfellow et al., 2014; Im et al., 2016; Smith & Topin, 2017; Li et al., 2018): for a 2-D landscape, we visualize the loss landscape along a specific direction $\bm{\delta}\in\mathbb{R}^{d}$, reducing the problem to visualizing a single-variable function: $L(\alpha)=\mathcal{T}\circ\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$. For a 3-D landscape, we visualize along two random directions $\bm{\delta}_{1},\bm{\delta}_{2}\in\mathbb{R}^{d}$ using $L(\alpha,\beta)=\mathcal{T}\circ\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta}_{1}+\beta\bm{\delta}_{2})$.
@@ 3.2 Most-case Loss Landscape
$L(\alpha)=\mathcal{T}\circ\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta}),\;\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I}).$ (1)
Empirically, we observe that different directions $\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})$ yield nearly identical results. Since $\mathcal{N}(\bm{0},\bm{I})$ represents uniformly random directions, this suggests that most directions produce similar landscapes. Therefore, we designate this as the most-case loss landscape. We only plot one random direction for each most-case landscape, while quantitatively validating the global prevalence of this geometry via hypothesis testing in Sec.˜D.4.
General Geometry. As shown in Fig.˜1, the most-case loss landscape for each capability resembles a basin, within which the models perform nearly identically, and outside of which they rapidly lose all capabilities (Peng et al., 2024). In particular, the pre-training stage creates a “basic capability basin” that endows the model with fundamental language comprehension and conversational abilities. Subsequent alignment stages sequentially establish specific capability basins (e.g., safety (Zou et al., 2023), math (Cobbe et al., 2021), coding (Mark Chen, 2021)) near this basic capability basin. More interestingly, this “basin” phenomenon emerges and becomes larger as the model size increases. For models like the Qwen-0.5B model, the loss landscape resembles the 0-1 loss landscape for small models in Li et al. (2018), which is continuous and smooth. When the models become bigger, the basins become more significant and easier to observe. Based on this observation, we argue that as long as subsequent benign fine-tuning remains within the basin of a specific capability, the parameters will remain within this basin and thus will not compromise those capabilities.
@@
Hypothesis Testing. A common concern is whether the loss landscape along several random directions can truly represent the geometry of a high-dimensional landscape with infinite possible directions. While we cannot test all infinite directions, we can statistically characterize the expected geometry of the "majority" of directions. In fact, if a property holds across several random directions, hypothesis testing enables us to assert, with an arbitrary type-I error, the percentage of directions that satisfy this property. For example, determining the percentage of directions in a $d$-dimensional space that form a strict basin is a hypothesis testing problem aimed at obtaining a statistical lower bound for the expectation of a Bernoulli variable $\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathbb{I}\{\mathcal{T}\circ\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0\}]$. We use the Clopper-Pearson bound here. Specifically, for Qwen2.5-7B on the AdvBench benchmark, we establish with 99% confidence that over 90% of all possible directions form a strict basin at a perturbation scale of $\sigma=0.01$. This hypothesis testing confirms that our "most-case" finding is a robust global property rather than a sampling artifact. See Sec.˜D.4 for detailed configurations and results.
The 0-1 Capability Landscape. Note that the basin phenomenon occurs primarily when using generative-based benchmarks. When using likelihood-based benchmarks, the loss landscape remains smooth and continuous. See Sec.˜C.1 for details. However, the loss landscape using benchmarks remains a valid loss landscape, since all benchmarks used in this paper employ a 0-1 loss on the dataset $\mathcal{D}$, where each sample is assigned a loss of 0/1 for a correct/incorrect response, which is widely studied (Keskar et al., 2017; Li et al., 2018; Garipov et al., 2018).
The Non-Trivial Nature of Basins. We argue that the observed basin structure is not merely a byproduct of probability thresholding. As detailed in Sec.˜D.9, we observe a region of semantic stability within the basin, where the model’s generated sentences often change structurally while the final answer remains correct (even under deterministic greedy decoding). We hypothesize that this phenomenon likely emerges from the combined effects of mode connectivity and the implicit bias of SGDs towards flatter minima (see Secs.˜5.3 and 7). See Sec.˜C.1 for a detailed discussion.
@@ 3.3 Worst-case Loss Landscape
The constraint $\|\bm{\delta}\|_{2}^{2}=\mathbb{E}[\|\mathcal{N}(\bm{0},\bm{I})\|_{2}^{2}]$ ensures that the perturbation norm matches that used in Figs.˜1 and 2, enabling more direct comparison for each model. Eq.˜2 is solved by optimizing $L(\bm{\theta}+\alpha\bm{\delta})$ using SGD and projecting the norm of $\bm{\delta}$ to unity at each step (Madry et al., 2018).
@@ 3.4 SFT-case Loss Landscape
Results. As shown in Fig.˜3(a), the loss landscape along the benign fine-tuning direction resembles the most-case landscape. It preserves safety within the most-case basin and loses capability when moving outside this basin. In Fig.˜3(b), when there is a distributional gap between the SFT dataset and the original dataset, the loss landscape becomes narrower and sharper, indicating that the fine-tuning direction does not align with the most-case directions. In Fig.˜3(c), when fine-tuned on the adversarial dataset, the model quickly learns to output “Sure, here is” as its initial tokens; consequently, its safety guardrails collapse rapidly, even though performance on other capability benchmarks remains intact. Thus, while some SFT configurations align closely with the most-case direction (e.g., Fig.˜3(a), math, general and code in Fig.˜3(c)), others deviate and cause capabilities to degrade more rapidly (e.g., Fig.˜3(b) and safety in Fig.˜3(c)). This variation clearly depends on the dataset and hyperparameters.
@@ 4 Theoretical Benefits of Basins
In this section, we adopt a soft definition of basins, similar to the flatness definition in Andriushchenko et al. (2023):
Definition 4.1.
A model $f_{\bm{\theta}}$ is said to have a $\sigma$-basin on benchmark $\mathcal{S}_{\mathcal{D}}$, if its noised version $f_{\bm{\theta}+\bm{\epsilon}}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$, performs nearly the same as the original version $f_{\bm{\theta}}$, i.e.,
$\mathcal{S}_{f,\mathcal{D}}(\bm{\theta})-\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\leq\tau.$ (3)
Definition 3 is a necessary condition for a model to have a most-case landscape resembling that in Fig.˜1: when $\tau\to 0$, it becomes the strict definition of basins, as defined by $\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathbb{I}\{\mathcal{T}\circ\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})=0\}]$ in Sec.˜3.2 (assuming performance does not increase under noise perturbation). Since our theoretical analysis holds for all $\tau$ rather than only $\tau=0$, we adopt this definition in this section to provide a theoretical analysis for a more general case.
@@ 4.1 Any Alignment Can Be Bounded by Average-case Alignment
This is achieved through the concept of randomized smoothing (Cohen et al., 2019; Salman et al., 2019; Lee et al., 2019). Since the model performs nearly identically within a $\sigma$-basin, for any input $\bm{x}$, instead of returning $f_{\bm{\theta}}(\bm{x})$, we can sample $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and return $f_{\bm{\theta}+\bm{\epsilon}}(\bm{x})$. Thus, the benchmark value for this model is $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})$. The following theorem demonstrates that, for any bounded benchmark $\mathcal{S}_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$ [^2] Without loss of generality, any benchmark with a bounded output range can be normalized to this interval to obtain a corresponding certified bound., regardless of how sensitive $f$ is to its parameters, the smoothed model $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ is at most $\frac{1}{\sqrt{2\pi}\sigma}$-Lipschitz. Consequently, when $\bm{\theta}_{0}$ is updated to $\bm{\theta}_{sft}$, the benchmark value changes by at most $\frac{1}{\sqrt{2\pi}\sigma}\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}$.
Theorem 4.2.
(Weak Law of Randomized Smoothing (Salman et al., 2019)) For any benchmark $\mathcal{S}_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, the function $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ is at most $\frac{1}{\sqrt{2\pi}\sigma}$-Lipschitz. Thus, we can bound the performance degradation as:
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\geq\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]-\frac{1}{\sqrt{2\pi}\sigma}\cdot\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}.$ (4)
@@
(Strong Law of Randomized Smoothing (Cohen et al., 2019; Salman et al., 2019; Lee et al., 2019; Chen et al., 2025a)) For any benchmark $\mathcal{S}_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, we have:
@@
$\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]\right)-\frac{\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}}{\sigma}\right),$ (5)
@@
where $\Phi$ is the cumulative distribution function of $\mathcal{N}(0,1)$, i.e., $\Phi(x)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{x}\exp\left(-\frac{s^{2}}{2}\right)ds$.
@@
The term $\mathbb{E}_{\bm{\epsilon}}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$ represents the expected benchmark value when sampling $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ and evaluating the benchmark on $f_{\bm{\theta}+\bm{\epsilon}}$. Eq.˜5 provides a guarantee of the expected benchmark value during fine-tuning. In practice, one typically samples a single $\bm{\epsilon}$ and evaluates $f_{\bm{\theta}+\bm{\epsilon}}$, rather than computing the expectation via extensive Monte Carlo sampling. Empirically, sampling a single instance yields results comparable to multiple samples. Theoretically, the variance introduced by sampling can be bounded using the well-known concentration phenomenon, which states that a random variable is unlikely to deviate significantly from its expectation:
Theorem 4.4.
@@
(Concentration of Gaussian, adapted from Wainwright (2019)) With probability at least $1-\delta$, we have $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})\geq\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]-\mathcal{L}\sigma\sqrt{2\log\frac{1}{\delta}}$, where $\mathcal{L}$ is the Lipschitz constant of $\mathcal{S}_{f,\mathcal{D}}$ with respect to $\bm{\theta}$.
Theoretical Positioning. We clarify our theoretical positioning here: while our mathematical derivations leverage these established techniques, our contribution lies in the fundamental conceptual shift of applying RS to the parameter space of LLMs. Unlike prior works that focus on input or data robustness, this novel application allows us to certify robustness against fine-tuning degradation, unifying benign and adversarial fine-tuning under a single framework. We provide a detailed discussion on this perspective and its reliance on basin stability in Sec.˜B.2.